find the x and y intercept of y=-9x-6
Answer by Gator Tutoring(5)

find the x and y intercept of y=-9x-6
To find the x intercept, make y equal to zero and solve for x.
0 = - 9x - 6
add 9x to both sides
9x = - 6
divide both sides by 9
x = - 6/9
simplify by dividing the top and bottom by 3
x = - 2/3 => x intercept
To find the y intercept, make x equal to zero and solve for y.
y = - 9 * 0 - 6
y = - 6 => y intercept
Gator Tutoring
pichonz@gmail.com
Getting Math help couldn't get any easier.
How?
1. Email me all your questions (minimum of 5)
2. I post worked out solutions on Google/Amazon/Ebay
3. You purchase, I send file through email.
Costs:
Standard = Worked out solutions
Premium = Worked out solutions + strategies + definitions + comments. Answers from above are done using the premium service.
Standard > $1 per problem
Premium > $2 per problem
note: Problems with multiple questions are counted as separate problems.